# Certifying rings of integers in number fields

### v1

@@ Abstract.
ccs: Security and privacy Logic and verificationccs: Mathematics of computing Mathematical software
@@ 1. Introduction
There are many fundamental concepts in mathematics that are, in principe, amenable to computation. Focussing on number theory, and more particularly algebraic number theory (i.e. the theory of algebraic numbers), such concepts include the rings of integers of a number field. Number fields generalize the field of rational numbers $\mathbf{Q}$ (in the sense that they are finite degree field extensions thereof) and each contains a ring of integers, which can be seen as generalizing how the (ordinary) integers $\mathbf{Z}$ are contained in the number field $\mathbf{Q}$; see Section 2 for more details. Rings of integers are thereby key to the arithmetic properties of their number fields, and both are completely essential to modern (and in fact much of 19th and 20th century) number theory. Their basic definitions and properties (amongst other concepts) were formalized in (Baanen et al., 2022), on which we build. From a computational direction, number fields of degree $2$ and their rings of integers were considered in (Baanen et al., 2023). This paper can be seen as a far reaching generalization of such formally verified computations.
The isomorphism class of a number field can be represented explicitly, for example by a polynomial with integer coefficients and leading coefficient $1$ that is irreducible over $\mathbf{Z}$ (and hence $\mathbf{Q}$). This is also a basic way to define number fields in Computer Algebra Systems such as PARI/GP (The PARI Group, 2023), SageMath (The Sage Developers, 2024), and Magma (Bosma et al., 1997) (perhaps relaxing to rational coefficients). Given the prime factorization of the discriminant of the defining polynomial, there exist efficient algorithms to determine the ring of integers, e.g. (Cohen, 1993, Algorithm 6.1.8), (Ford, 1978, Chapter III), and implementations are available in various Computer Algebra Systems, including those mentioned before. However, these algorithms are quite involved. In order to work with rings of integers inside a proof assistant —Lean 4 in our case— we therefore opted for a certification approach. Given a basis for the ring of integers, e.g. computed by an external Computer Algebra System or extracted from some database, we let SageMath compute a certificate which is checked in Lean to certify the correctness of the ring of integers. In fact, our SageMath code outputs a complete proof which Lean can check immediately. This is described in Section 4, the technical heart of this paper.
@@
In order for everything to function, we needed to formalize much underlying theory and computational methods, most of which we believe is also interesting in its own rights. Some notable topics include irreducibility of polynomials over $\mathbf{Q}$ and finite fields (see Section 3), as well as resultants and discriminants (see Section 5).
@@ 2. Preliminaries
In this paper, we assume some familiarity with basic ring and field theory, as can be found e.g. in the undergraduate textbooks (Dummit and Foote, 2004; Lang, 2005). For sake of self containedness, we will discuss in this section some basics concerning number fields and their rings of integers; the following paragraph essentially follows the exposition (with very little modification) from (Baanen et al., 2023, Section 2).
@@
The last (cubic) example shows that the ring of integers can become complicated very quickly. In this case $\mathcal{O}_{\mathbf{Q}(\alpha)}$ cannot be written in the form $\mathbf{Z}[\gamma]$ for any element $\gamma\in\mathcal{O}_{\mathbf{Q}(\alpha)}$, which is phrased as $\mathcal{O}_{\mathbf{Q}(\alpha)}$ being not monogenic. Note that any number field (say of degree $n$) is monogenic, as it can be generated by a single element $\beta$ over $\mathbf{Q}$, which yields $\{1,\beta,\beta^{2},\ldots,\beta^{n-1}\}$ as a basis for the corresponding $\mathbf{Q}$-vector space, called a power basis. Note that the ring of integers $\mathcal{O}_{K}$ of a number field $K$ is a $\mathbf{Z}$-module. Recall that in general an $R$-module is just a vector space if $R$ is a (skew) field, but the same standard defining axioms make sense in the more general case that the scalars $R$ are only assumed to form a ring. While the $\mathbf{Z}$-module $\mathcal{O}_{K}$ might not have a power basis (as it may not be monogenic), it always has some basis, called a $\mathbf{Z}$-basis or integral basis. E.g. in the cubic example, a $\mathbf{Z}$-basis is given by $\{1,\alpha,(\alpha-\alpha^{2})/2\}$, as any element of $\mathcal{O}_{\mathbf{Q}(\alpha)}$ can be expressed uniquely as a $\mathbf{Z}$-linear combination of these three basis elements. Finally, if we have a subring $\mathcal{O}$ of a number field $K$, which as a $\mathbf{Z}$-module has a basis that also forms a $\mathbf{Q}$-basis for $K$, then $\mathcal{O}$ is actually contained in $\mathcal{O}_{K}$.
@@ 3. Irreducibility
For a primitive polynomial $f$ in $\mathbf{Z}[X]$, proving that $\mathbf{Q}[X]/\langle f\rangle$ is a number field is equivalent to proving that $f$ is irreducible over $\mathbf{Z}[X]$. Thus, generating an irreducibility proof for polynomials over the integers in Lean is crucial for our purposes.
@@ 3.1. Polynomials as lists
Additionally, we defined ComputablePolynomial R as the subtype of lists $l$ over R such that l = l.dropTrailingZeros, and proved that it is isomorphic to Mathlib polynomials. This representation is similar to the one in Coq’s Mathematical Components library (Mahboubi and Tassi, 2020), which uses lists with a non-zero last entry to define polynomials, though our approach does not assume that $R$ is nontrivial.
@@ 3.2. Over finite fields
Let us start by considering irreducibility of polynomials over a finite field, which will help us prove irreducibility in $\mathbf{Z}[X]$.
@@
In Lean, to talk about a finite field, we endow (F : Type*) with instances [Field F] [Fintype F]. The previous statement –which we formalized– reads:
@@
Arithmetic in finite fields has only been implemented in Lean for fields of prime order (our work on irreducible polynomials can contribute to extending this to higher degree extensions), therefore we focus on the case $q=p$, with $p$ a prime number. In Lean, the integers modulo $p$ are represented as ZMod p, which has decidable equality, allowing us to solve identity goals simply with the decide tactic. Furthermore, if we have the instance Fact (Nat.Prime p) available, then Lean automatically infers that ZMod p is a field.
Rabin’s test starts by proving $f\mid X^{p^{n}}-X$. The obvious way to do this is to provide the factor $s$ such that $X^{p^{n}}-X=f*s$ and verify the equality. There is, however, a problem. In general, $s$ will not be sparse and it will be of huge degree equal to $p^{n}-n$. This makes storing it and verifying the corresponding identity infeasible. The standard approach taken in computer algebra  (von zur Gathen and Gerhard, 2013) is instead to show that $X^{p^{n}}\equiv X\pmod{f}$ using a square-and-multiply algorithm, reducing modulo $f$ at each step by performing polynomial division with remainder. We currently do not have an effective implementation of polynomial division with remainder in Lean. However, with our implementation of polynomials as lists, we can quickly verify multiplications. This suggest a certificate for $f\mid X^{p^{n}}-X$ based on the square-and-multiply algorithm.
This well-known algorithm performs fast exponentiation on any monoid. When applied to the ring $\mathbf{F}_{p}[X]/\langle f\rangle$, the underlying proposition looks like:
@@
Let $g$ in $\mathbf{F}_{p}[X]$ and let $n$ be a non-negative integer. Let $(b_{i})_{i=0}^{s}$ be the binary digits of $n$ so that $n=\sum_{i=0}^{s}b_{i}2^{i}$. Set $y_{s}\equiv g^{b_{s}}\pmod{f}$ and $y_{i}\equiv y_{i+1}^{2}*g^{b_{i}}\pmod{f}$ for $i=s-1,\ldots,0$. Then $y_{0}\equiv g^{n}\pmod{f}$.
Computing $g^{n}\pmod{f}$ this way requires significantly fewer operations than repeated multiplication. The polynomials $y_{i}$ can be chosen with degrees less than $n$. Note that, when $a$ and $b$ do not have very large degree, we can certify $a\equiv b\pmod{f}$ in the obvious way.
@@
Let $f$ in $\mathbf{F}_{p}[X]$ of degree $n>0$, let $(b_{i})_{i=0}^{s}$ be such that $p=\sum_{i=0}^{s}b_{i}2^{i}$. A certificate for the irreducibility of $f$ consists of the following data:
@@
    $a_{n/t}*f+b_{n/t}*(h_{n/t}-X)=1$ for every prime $t\mid n$
The first two statements in the verification imply that $h_{i}^{p}\equiv h_{i+1}\pmod{f}$ for all $0\leq i<n$. Together with $h_{0}=X$, we get that $X^{p^{i}}\equiv h_{i}\pmod{f}$ for every $0\leq i\leq n$. Since $h_{n}=X$, the first part of Rabin’s test is shown. The fourth statement proves the second part of the test.
@@
For an irreducible polynomial over $\mathbf{F}_{p}$, a certificate of this form always exists.The function CertificateIrrModpFFP, implemented in SageMath, computes the certifying data and writes the corresponding term in Lean. For small degrees (usually $n\leq 12$) all of the proofs included as fields in this structure can be solved automatically either by using the rfl or the decide tactic. For higher degrees, the computation may exceed Lean’s preset maximum recursion depth or heartbeats limits.
@@ 3.3. Degree analysis
Proving irreducibility over $\mathbf{F}_{p}[X]$ provides us with a way to demonstrate irreducibility of many polynomials in $\mathbf{Z}[X]$.
Let $f$ in $\mathbf{Z}[X]$ of degree $n$ and $p$ a prime number not dividing $\operatorname{lc}(f)$, the leading coefficient of $f$. Let $\bar{f}=f\pmod{p}$ be the polynomial in $\mathbf{F}_{p}$ obtained by reducing the coefficients of $f$ modulo $p$. Write $\bar{f}=\prod_{i=1}^{m}g_{i}$, with $g_{i}$ irreducible in $\mathbf{F}_{p}[X]$. Consider the tuple $D_{p}=(\deg(g_{i}))_{i}$ and let $S(D_{p})$ be the set of all possible subset sums of $D_{p}$. It is not hard to see that if $a$ divides $f$, then $\deg a\in S(D_{p})$. Computing $S(D_{p})$ for different primes $p$ will thus give us refined information on the degree of the factors of $f$. In particular:
@@
Let $f$ in $\mathbf{Z}[X]$ be primitive and $p_{1},\ldots,p_{m}$ a set of distinct primes not dividing $\operatorname{lc}(f)$. Let

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 14th ACM SIGPLAN International Conference on Certified Programs and Proofs; January 20–21, 2025; Denver, CO, USAbooktitle: Proceedings of the 14th ACM SIGPLAN International Conference on Certified Programs and Proofs (CPP ’25), January 20–21, 2025, Denver, CO, USAdoi: 10.1145/3703595.3705874isbn: 979-8-4007-1347-7/25/01ccs: Security and privacy Logic and verificationccs: Mathematics of computing Mathematical software
@@ 1. Introduction
There are many fundamental concepts in mathematics that are, in principle, amenable to computation. Focusing on number theory, and more particularly algebraic number theory (i.e. the theory of algebraic numbers), such concepts include the rings of integers of a number field. Number fields generalize the field of rational numbers $\mathbf{Q}$ (in the sense that they are finite degree field extensions thereof) and each contains a ring of integers, which can be seen as generalizing how the (ordinary) integers $\mathbf{Z}$ are contained in the number field $\mathbf{Q}$; see Section 2 for more details. Rings of integers are thereby key to the arithmetic properties of their number fields, and both are essential to number theory from the 19th century to the modern day. Their basic definitions and properties (amongst other concepts) were formalized in (Baanen et al., 2022), on which we build. From a computational direction, number fields of degree $2$ and their rings of integers were considered in (Baanen et al., 2023). This paper can be seen as a far reaching generalization of such formally verified computations. In order to work with rings of integers inside a proof assistant —Lean 4 in our case— we opted for a certification approach. Given a ring of integers, concretely represented by giving an integral basis (i.e. a $\mathbf{Z}$-basis) for it, as e.g. computed by an external Computer Algebra System (CAS) or extracted from some database, we let SageMath (The Sage Developers, 2024) compute a certificate which is checked in Lean to certify the correctness of the ring of integers. In fact, our SageMath code outputs a complete proof which Lean can readily check.
The isomorphism class of a number field can be represented explicitly, for example by a polynomial with integer coefficients and leading coefficient $1$ that is irreducible over $\mathbf{Z}$ (and hence $\mathbf{Q}$). This is also a basic way to define number fields in Computer Algebra Systems such as PARI/GP (The PARI Group, 2023), SageMath (The Sage Developers, 2024), and Magma (Bosma et al., 1997) (perhaps relaxing to rational coefficients). Given the prime factorization of the discriminant of the defining polynomial, there exist efficient algorithms to determine the ring of integers, e.g. (Cohen, 1993, Algorithm 6.1.8), (Ford, 1978, Chapter III), and implementations are available in various Computer Algebra Systems, including those mentioned before. However, these algorithms are quite involved, which is one reason why we opted for a certification approach for rings of integers. This is described in Section 5, the technical heart of this paper.
@@
In Section 3, we provide an overview of the different notions of Lean computation used. In order for everything to function, we needed to formalize much underlying theory and computational methods, most of which are interesting independently. Notable topics include irreducibility of polynomials over $\mathbf{Q}$ and finite fields (see Section 4), as well as resultants and discriminants (see Section 6).
@@ 2. Preliminaries
In this paper, we assume some familiarity with basic ring and field theory, as can be found e.g. in the undergraduate textbooks (Dummit and Foote, 2004; Lang, 2005). For sake of self-containedness, we will discuss in this section some basics concerning number fields and their rings of integers; the following paragraph essentially follows the exposition (with very little modification) from (Baanen et al., 2023, Section 2).
@@
The last (cubic) example shows that the ring of integers can become complicated very quickly. In this case $\mathcal{O}_{\mathbf{Q}(\alpha)}$ cannot be written in the form $\mathbf{Z}[\gamma]$ for any element $\gamma\in\mathcal{O}_{\mathbf{Q}(\alpha)}$, which is phrased as $\mathcal{O}_{\mathbf{Q}(\alpha)}$ being not monogenic. Note that any number field (say of degree $n$) is monogenic, as it can be generated by a single element $\beta$ over $\mathbf{Q}$, which yields $\{1,\beta,\beta^{2},\ldots,\beta^{n-1}\}$ as a basis for the corresponding $\mathbf{Q}$-vector space, called a power basis. Observe that the ring of integers $\mathcal{O}_{K}$ of a number field $K$ is a $\mathbf{Z}$-module. Recall that in general an $R$-module is just a vector space if $R$ is a (skew) field, but the same standard defining axioms make sense in the more general case that the scalars $R$ are only assumed to form a ring. While the $\mathbf{Z}$-module $\mathcal{O}_{K}$ might not have a power basis (as it may not be monogenic), it always has some basis, called a $\mathbf{Z}$-basis or integral basis. E.g. in the cubic example, a $\mathbf{Z}$-basis is given by $\{1,\alpha,(\alpha-\alpha^{2})/2\}$, as any element of $\mathcal{O}_{\mathbf{Q}(\alpha)}$ can be expressed uniquely as a $\mathbf{Z}$-linear combination of these three basis elements. Finally, if we have a subring $\mathcal{O}$ of a number field $K$, which as a $\mathbf{Z}$-module has a basis that also forms a $\mathbf{Q}$-basis for $K$, then $\mathcal{O}$ is actually contained in $\mathcal{O}_{K}$.
3. Forms of Computation in Lean
This section provides an overview of the different notions of computation that we had to deal with.
Lean has a built-in notion of computation through reduction: this is a primitive relation between two terms reflecting the computational content of the Calculus of Inductive Constructions (Carneiro, 2019). Reduction is invoked as part of definitional equality checking, and allows Lean to verify equalities such as $2+2=1+3$ through computing that either side evaluates to $4$. Definitionally equal terms are indistinguishable for the type theory, and therefore a proof of $2+2=1+3$ can be given by the reflexivity principle of equality, rfl. In Lean 4, reduction also applies to  literals of primitive types: natural numbers and strings. Reduction is implemented as part of the trusted codebase in the Lean kernel.
Lean as a programming language additionally possesses a notion of computation for definitions which can be compiled into executable code and evaluated. Evaluation is somewhat orthogonal to reduction in that it is not part of the logic of Lean, and not all definitions that can be evaluated will reduce, and vice versa. In particular, Lean includes –and Mathlib makes consistent use of– the axiom of choice, resulting in noncomputable definitions that cannot be evaluated. The user can additionally override with arbitrary code the computational meaning of a definition, using the @[implemented_by] attribute.
Evaluation is the notion of computation powering the tactic framework. Broadly speaking, tactics are evaluated Lean code that operate directly on the term level. Tactics can also be said to perform computation, by taking in a term t and producing a new term t alongside a proof term of type t = t. The norm_num tactic for computing with numerical expressions is an example of this design. Often, a tactic works by reflecting a given term into a data structure internal to the tactic, using computation on internal data to construct a proof term. Since tactics can choose the internal representation of the data structures, they reimbue noncomputable terms with computational meaning. An example is the ring tactic of Mathlib which computes on polynomial expressions although the ring structure of the type Polynomial itself is marked noncomputable.
If a tactic returns rfl as the proof term for the equality t = t, verifying the definitional equality of the terms t and t can invoke reduction: this principle of computational reflection (Boutin, 1997) saves the need to go through the computation twice, once on the tactic-internal data structure and once on the constructed proof term. The tactic decide is an example of a reflective tactic in Lean, proving a given goal P by finding a corresponding decision procedure d such that d = true if and only if P, and then verifying that d = true holds. This tactic has two variants: decide first computes whether d = true in the elaborator, before submitting the same computation to the kernel, while decide! directly invokes kernel reduction and thus can unfold every definition.
Efficiency is an important consideration in the use of computation in Lean. In the highest generality, evaluation is faster than reduction since evaluation does not rely on the abstraction of terms that must be constructed and matched for each reduction step. The higher efficiency of evaluation comes at the cost of being separated from the logic: proof checking can only invoke reduction. Lean provides an escape hatch in the form of the native_decide tactic: unlike decide, this allows a proof whenever a decision procedure evaluates to true. Use of native_decide not only means that the entire evaluation mechanism has to be added to the trusted codebase, in addition the correctness of all @[implemented_by] attributes must be trusted. Use of native_decide is therefore a tradeoff between efficiency and certainty. See Section 7 for some performance measurements in our project.
@@ 4. Irreducibility
For a primitive polynomial $f$ in $\mathbf{Z}[X]$ – a polynomial whose coefficients have greatest common divisor equal to $1$– proving that $\mathbf{Q}[X]/\langle f\rangle$ is a number field is equivalent to proving that $f$ is irreducible over $\mathbf{Z}[X]$. Thus, generating an irreducibility proof for polynomials over the integers in Lean is crucial for our purposes.
@@ 4.1. Polynomials as Lists
We note that this approach to polynomial computation relies on the input polynomials being provided in the list-based format, as the standard way to define polynomials in Mathlib is noncomputable. Alternatively, one could consider using the Mathlib representation of polynomials along with tactics such as ring (Baanen, 2020) (which is partly based on the Coq tactic (Grégoire and Mahboubi, 2005)) and norm_num extensions like reduce_mod_char to prove equalities involving additions and multiplications. However, this method proved too slow for our purposes, which led us to adopt the list-based approach instead.
In addition to Polynomial.ofList and the arithmetic operations on lists, we defined ComputablePolynomial R as the subtype of lists $l$ over R such that l = l.dropTrailingZeros, and proved that it is isomorphic to Mathlib polynomials. The algebraic structure of ComputablePolynomial R, namely, that it is a ring when R is a ring, is pulled from the corresponding instance in Mathlib polynomials. This representation is similar to the one in Coq’s Mathematical Components library (Mahboubi and Tassi, 2020), which uses lists with a non-zero last entry to define polynomials, though our approach does not assume that $R$ is nontrivial.
Having multiple encodings for the same mathematical object, one better suited for proving theoretical properties and another for computation, has been explored in previous formalization efforts of algebraic algorithms (Dénès et al., 2012). While we do not employ a proof transfer framework such as Trocq (Cohen et al., 2024), this could potentially be explored to automate translations between equivalent data structures.
@@ 4.2. Over Finite Fields
In this section, we describe a way to prove irreducibility of polynomials over a finite field, which will aid us in proving irreducibility in $\mathbf{Z}[X]$. We then present a certificate for this irreducibility and its implementation in Lean.
@@
In Lean, to talk about a finite field, we follow the practice in Mathlib and endow (F : Type*) with instances [Field F] [Fintype F]. The instance [Fintype F] carries data: a finite set containing all elements of F. We use it rather than [Finite F], which merely asserts that F is in bijection with Fin n, the canonical type with n elements, for some natural number n. The previous statement –which we formalized– reads:
@@
The term Fintype.card F is the cardinality of the finite set of all elements of F coming from the [Fintype F] instance.
Currently, automatic computations in finite fields can only be carried out in Lean for fields of prime order (our work on irreducible polynomials can contribute to extending this to higher degree extensions), therefore we focus on the case $q=p$, with $p$ a prime number. In Lean, the integers modulo $p$ are represented as ZMod p, which has decidable equality, allowing us to solve identity goals simply with the decide tactic. Furthermore, if we have the instance Fact (Nat.Prime p) available, then Lean infers that ZMod p is a field.
Rabin’s test starts by proving $f\mid X^{p^{n}}-X$, where $n$ is the degree of the polynomial $f$ in $\mathbf{F}_{p}[X]$. The obvious way to do this is to provide the factor $s$ such that $X^{p^{n}}-X=f*s$ and verify the equality. There is, however, a problem. In general, $s$ will not be sparse and it will be of huge degree equal to $p^{n}-n$. This makes storing it and verifying the corresponding identity infeasible. The standard approach taken in computer algebra (von zur Gathen and Gerhard, 2013) is instead to show that $X^{p^{n}}\equiv X\pmod{f}$ using a square-and-multiply algorithm, reducing modulo $f$ at each step by performing polynomial division with remainder. We currently do not have an effective implementation of polynomial division with remainder in Lean. However, with our implementation of polynomials as lists, we can quickly verify multiplications. This suggest a certificate for $f\mid X^{p^{n}}-X$ based on the square-and-multiply algorithm.
This well-known algorithm performs fast exponentiation on any monoid. When applied to the ring $\mathbf{F}_{p}[X]/\langle f\rangle$, the underlying proposition looks like this:
@@
Let $f$ and $g$ in $\mathbf{F}_{p}[X]$ and let $m$ be a non-negative integer. Let $(\beta_{i})_{i=0}^{s}$ be the binary digits of $m$ so that $m=\sum_{i=0}^{s}\beta_{i}2^{i}$. Set $y_{s}\equiv g^{\beta_{s}}\pmod{f}$ and $y_{i}\equiv y_{i+1}^{2}*g^{\beta_{i}}\pmod{f}$ for $i=s-1,\ldots,0$. Then $y_{0}\equiv g^{m}\pmod{f}$.
Computing $g^{m}\pmod{f}$ this way requires significantly fewer operations than repeated multiplication. The polynomials $y_{i}$ can be chosen with degrees less than $n$. Note that, when $a$ and $b$ do not have very large degree, we can certify $a\equiv b\pmod{f}$ in the obvious way.
@@
Let $f$ in $\mathbf{F}_{p}[X]$ of degree $n>0$, let $(\beta_{i})_{i=0}^{s}$ be the binary digits of $p$, so $p=\sum_{i=0}^{s}\beta_{i}2^{i}$. A certificate for the irreducibility of $f$ consists of the following data:
@@
    $a_{n/t}*f+b_{n/t}*(h_{n/t}-X)=1$ for every prime $t\mid n$.
The first two statements in the verification imply that $h_{i}^{p}\equiv h_{i+1}\pmod{f}$ for all $0\leq i<n$. Together with $h_{0}=X$, we get that $X^{p^{i}}\equiv h_{i}\pmod{f}$ for every $0\leq i\leq n$. Since $h_{n}=X$, the first part of Rabin’s test is shown. The fourth statement proves the second part of the test. Note that in this last statement, only some entries $a_{i}$ and $b_{i}$, with $i$ dividing $n$, are used. Consequently, the rest of the entries in these $n$-tuples, which are included solely to simplify the indexing, can be set to zero.
@@
For an irreducible polynomial over $\mathbf{F}_{p}$, a certificate of this form always exists. The function CertificateIrrModpFFP, implemented in SageMath, computes the certifying data –including the decomposition of the degree into prime factors– and writes the corresponding term in Lean. For small degrees (usually $n\leq 12$) all of the proofs included as fields in this structure can be solved automatically either by using the rfl or the decide tactic. For higher degrees, the computation may exceed Lean’s preset maximum recursion depth or heartbeats limits. A variety of approaches could be used to deal with this timeout issue: it might be possible to develop better decision procedures for use with decide, or we might use (a custom extension to) the norm_num tactic that performs this computation.
@@ 4.3. Degree Analysis
Here, we describe a certificate for the irreducibility of polynomials in $\mathbf{Z}[X]$, constructed by combining information from the factorization of the polynomial modulo distinct primes.
Let $f$ in $\mathbf{Z}[X]$ of degree $n$ and $p$ a prime number not dividing $\operatorname{lc}(f)$, the leading coefficient of $f$. Let $\bar{f}=f\pmod{p}$ be the polynomial over $\mathbf{F}_{p}$ obtained by reducing the coefficients of $f$ modulo $p$. Write $\bar{f}=\prod_{i=1}^{m}g_{i}$, with $g_{i}$ irreducible in $\mathbf{F}_{p}[X]$. Consider the tuple $D_{p}=(\deg(g_{i}))_{i}$ and let $S(D_{p})$ be the set of all possible subset sums of $D_{p}$. It is not hard to see that if $a$ divides $f$, then $\deg a\in S(D_{p})$. Computing $S(D_{p})$ for different primes $p$ will thus give us refined information on the degree of the factors of $f$. In particular:
@@
Let $f$ in $\mathbf{Z}[X]$ be primitive and $p_{1},\ldots,p_{m}$ a collection of distinct primes not dividing $\operatorname{lc}(f)$. Let